# Stochastic Momentum Methods for Non-smooth Non-Convex Finite-Sum Coupled Compositional Optimization

### v1

@@ Appendix D Convergence Analysis of ALEXR2 applying to Constraint Optimization Problems
$+\left(\frac{1}{2\eta}+\frac{\rho_{g}}{2}\right)\mathbb{E}\left\|\mathbf{z}_{*}-\mathbf{z}_{k}\right\|_{2}^{2}-\left(\frac{1}{2\eta}+\frac{1}{4\nu}\right)\mathbb{E}\left\|\mathbf{z}_{*}-\mathbf{z}_{k+1}\right\|_{2}^{2}$
$-\left(\frac{1}{m\gamma}-\frac{\lambda_{2}+\lambda_{3}\theta}{\lambda m}\right)\mathbb{E}\left[U_{\bar{f}_{i}^{*}}(\bar{\mathbf{y}}_{k+1},\mathbf{y}_{k})\right]-\left(\frac{1}{2\eta}-\frac{C_{g}^{2}}{2\lambda_{2}}-{\frac{\frac{1}{\eta}+\frac{1}{\nu}}{4}}\right)\mathbb{E}\left\|\mathbf{z}_{k+1}-\mathbf{z}_{k}\right\|_{2}^{2}$
$+\frac{\theta C_{g}^{2}}{2\lambda_{3}}\mathbb{E}\left\|\mathbf{z}_{k}-\mathbf{z}_{k-1}\right\|_{2}^{2}+\mathbb{E}[\Gamma_{k+1}-\theta\Gamma_{k}]+\frac{2(1+2\theta)\sigma_{0}^{2}}{B_{2}\lambda(1+\gamma^{-1})}+\frac{\frac{C_{f}^{2}\sigma_{1}^{2}}{B_{2}}+\frac{\sigma^{2}}{B_{1}}+8C_{f}^{2}C_{g}^{2}+8C_{g}^{2}}{\frac{1}{\eta}+\frac{1}{\nu}}.$ (45)
@@ Appendix E SONEX with Adaptive Learning Rates
$\mathbf{w}_{t+1}=\mathbf{w}_{t}-\tilde{\eta}\circ\mathbf{v}_{t+1},\tilde{\eta}=\frac{\eta}{\sqrt{\mathbf{s}_{t}}+\varepsilon},\mathbf{s}_{t+1}=(1-\beta^{\prime})\mathbf{s}_{t}+\beta^{\prime}G_{t}^{2}$
where $G_{t}$ is the overall gradient estimator mentioned in algorithm 1 and $G_{t}^{2}$ denote an element-wise square of $G_{t}$. The following assumption has been justified for the adaptive step size by [51].
@@
Note that $\left\|\mathbf{w}_{t+1}-\mathbf{w}_{t}\right\|^{2}\leq\left\|\tilde{\eta}\circ\mathbf{v}_{t+1}\right\|^{2}\leq\left\|\eta^{2}c_{u}^{2}\mathbf{v}_{t+1}\right\|^{2}=\eta^{2}c_{u}^{2}\left\|\mathbf{v}_{t+1}\right\|^{2}$, we can easily get a similar bound for $V_{t+1},U_{t+1}$ as follow:
@@ F.2 AUC Maximization with ROC Fairness Constraints
$h_{\tau}^{+}(\mathbf{w}):=\Big{|}\frac{1}{n_{p}^{+}}\sum_{i=1}^{n_{p}}\mathbf{I}\{b^{p}_{i}=1\}\sigma(s_{\mathbf{w}}(\mathbf{a}_{i}^{p})-\tau)\vspace*{-0.1in}-\frac{1}{n_{u}^{+}}\sum_{i=1}^{n_{u}}\mathbf{I}\{b^{u}_{i}=1\}\sigma(s_{\mathbf{w}}(\mathbf{a}_{i}^{u})-\tau)\Big{|}-\kappa\leq 0$
$h_{\tau}^{-}(\mathbf{w}):=\Big{|}\frac{1}{n_{p}^{-}}\sum_{i=1}^{n_{p}}\mathbf{I}\{b^{p}_{i}=-1\}\sigma(s_{\mathbf{w}}(\mathbf{a}_{i}^{p})-\tau)\vspace*{-0.1in}-\frac{1}{n_{u}^{-}}\sum_{i=1}^{n_{u}}\mathbf{I}\{b^{u}_{i}=-1\}\sigma(s_{\mathbf{w}}(\mathbf{a}_{i}^{u})-\tau)\Big{|}-\kappa\leq 0,$
@@
It is notable that the penalty terms in the above problem can be formulated as $[h_{\tau}^{+}(\mathbf{w})]_{+}=f(g(\mathbf{w};\tau))$, where $f(g)=[|g_{1}-g_{2}|-\kappa]_{+}$, $g_{1}(\mathbf{w};\tau)=\frac{1}{n_{p}^{+}}\sum_{i=1}^{n_{p}}\mathbf{I}\{b_{i}^{p}=1\}\sigma(s_{\mathbf{w}}(\mathbf{a}_{i}^{p})-\tau)$, $g_{2}(\mathbf{w};\tau)=\frac{1}{n_{u}^{+}}\sum_{i=1}^{n_{u}}\mathbf{I}\{b_{i}^{u}=1\}\sigma(s_{\mathbf{w}}(\mathbf{a}_{i}^{u})-\tau)$. As a result, $f$ is convex, $g(\mathbf{w})$ is smooth. As a result, Algorithm 2 is applicable.
@@ F.3 Continual learning with non-forgetting constraints
[…]
@@
F.4 Verification of Assumption 4.3
@@
F.5 Other Details of Experiments

### v2

@@ Appendix D Convergence Analysis of SONEX and ALEXR2 applying to Constraint Optimization Problems
$+\left(\frac{1}{2\eta}+\frac{\sqrt{d_{1}}C_{f}\rho_{g}}{2}+\frac{\rho_{g}}{2}\right)\mathbb{E}\left\|\mathbf{z}_{*}-\mathbf{z}_{k}\right\|_{2}^{2}-\left(\frac{1}{2\eta}+\frac{1}{4\nu}\right)\mathbb{E}\left\|\mathbf{z}_{*}-\mathbf{z}_{k+1}\right\|_{2}^{2}$
$-\left(\frac{1}{m\gamma}-\frac{\lambda_{2}+\lambda_{3}\theta}{\lambda m}\right)\mathbb{E}\left[U_{\bar{f}_{i}^{*}}(\bar{\mathbf{y}}_{k+1},\mathbf{y}_{k})\right]-\left(\frac{1}{2\eta}-\frac{C_{g}^{2}}{2\lambda_{2}}-{\frac{3}{8}(\frac{1}{\eta}+\frac{1}{\nu})}\right)\mathbb{E}\left\|\mathbf{z}_{k+1}-\mathbf{z}_{k}\right\|_{2}^{2}$
$+\frac{\theta C_{g}^{2}}{2\lambda_{3}}\mathbb{E}\left\|\mathbf{z}_{k}-\mathbf{z}_{k-1}\right\|_{2}^{2}+\mathbb{E}[\Gamma_{k+1}-\theta\Gamma_{k}]+\frac{2(1+2\theta)\sigma_{0}^{2}}{B_{2}\lambda(1+\gamma^{-1})}+\frac{\frac{C_{f}^{2}\sigma_{1}^{2}}{B_{2}}+\frac{C_{f}^{2}C_{g}^{2}}{B_{1}}+8C_{f}^{2}C_{g}^{2}+8C_{g}^{2}}{\frac{1}{\eta}+\frac{1}{\nu}}.$ (44)
@@ Appendix E SONEX with Adaptive Learning Rates
$\mathbf{w}_{t+1}=\mathbf{w}_{t}-\tilde{\eta}\circ\mathbf{v}_{t+1},\tilde{\eta}=\frac{\eta}{\sqrt{\mathbf{s}_{t}}+\varepsilon},\mathbf{s}_{t+1}=(1-\beta^{\prime})\mathbf{s}_{t}+\beta^{\prime}G_{t}\circ G_{t}$
where $G_{t}$ is the overall gradient estimator mentioned in algorithm 1 and $\circ$ denotes Hadamard(element-wise) product. The following assumption has been justified for the adaptive step size by [48].
@@ F.1 GDRO with CVaR divergence
Motivation of GDRO Modern machine learning models are typically trained under the empirical risk minimization (ERM) framework, which treats all samples in the dataset equally. Although ERM often achieves strong average performance on test sets drawn from distributions similar to the training data, it can perform poorly on rare or underrepresented subpopulations, i.e., it lacks robustness to distributional shifts. The motivation of GDRO is to address the spurious correlation between features and labels, and distributional shift. GDRO partitions the dataset into groups representing different distributions and applies a robust optimization scheme that assigns more weight on the worst groups. Specifically, GDRO considers the following objective:
$\min_{\theta}\max_{\mathbf{p}\in\Omega}\frac{1}{n}\sum_{g=1}^{n}p_{g}\mathbb{E}_{(\mathbf{x},y)\sim\mathcal{D}_{g}}l(\theta;(\mathbf{x},y))$
where $\Omega\subseteq\Delta$ is the set of distribution under consideration and $\Delta$ is the simplex. We consider here a popular choice of $\Omega=\{\mathbf{p}\in\Delta:p_{i}\leq\frac{1}{k},\forall i\in[n]\}$, which, by solving the inner-level maximization problem, leads to an equivalent reformulation 10 corresponding to the so-called CVaR divergence [5, 49]. Intuitively GDRO with CVaR divergence aims to minimize averaged loss of the top-k worst groups.
@@ F.2 AUC Maximization with ROC Fairness Constraints
$h_{\tau}^{+}(\mathbf{w})=\Big|\frac{1}{n_{p}^{+}}\sum_{i=1}^{n_{p}}\mathbf{I}\{b^{p}_{i}=1\}\sigma(s_{\mathbf{w}}(\mathbf{a}_{i}^{p})-\tau)\vskip-7.22743pt-\frac{1}{n_{u}^{+}}\sum_{i=1}^{n_{u}}\mathbf{I}\{b^{u}_{i}=1\}\sigma(s_{\mathbf{w}}(\mathbf{a}_{i}^{u})-\tau)\Big|-\kappa\leq 0$
$h_{\tau}^{-}(\mathbf{w})=\Big|\frac{1}{n_{p}^{-}}\sum_{i=1}^{n_{p}}\mathbf{I}\{b^{p}_{i}=-1\}\sigma(s_{\mathbf{w}}(\mathbf{a}_{i}^{p})-\tau)\vskip-7.22743pt-\frac{1}{n_{u}^{-}}\sum_{i=1}^{n_{u}}\mathbf{I}\{b^{u}_{i}=-1\}\sigma(s_{\mathbf{w}}(\mathbf{a}_{i}^{u})-\tau)\Big|-\kappa\leq 0,$
@@
It is notable that the penalty terms in the above problem can be formulated as $h_{\tau}^{+}(\mathbf{w})]_{+}=f(g(\mathbf{w};\tau))$, where $f(g)=[|g_{1}-g_{2}|-\kappa]_{+}$, $g_{1}(\mathbf{w};\tau)=\frac{1}{n_{p}^{+}}\sum_{i=1}^{n_{p}}\mathbf{I}\{b_{i}^{p}=1\}\sigma(s_{\mathbf{w}}(\mathbf{a}_{i}^{p})-\tau)$, $g_{2}(\mathbf{w};\tau)=\frac{1}{n_{u}^{+}}\sum_{i=1}^{n_{u}}\mathbf{I}\{b_{i}^{u}=1\}\sigma(s_{\mathbf{w}}(\mathbf{a}_{i}^{u})-\tau)$. As a result, $f$ is convex, $g(\mathbf{w})$ is smooth. As a result, Algorithm 2 is applicable.
@@ F.3 Continual learning with non-forgetting constraints
$\text{s.t. }h_{k}:=\mathcal{L}_{k}(\mathbf{w},D_{k})-\mathcal{L}_{k}(\mathbf{w}_{\text{old}},D_{k})\leq 0,k=1,\cdots,m.$
@@
F.4 Ablation Study
We conduct a series of ablation studies to examine the effects of key hyperparameters and design choices.
Results for varying $\beta$
We investigate the impact of different $\beta$ on model performance. We evaluate $\beta\in\{0.1,0.3,0.5,0.7,1.0\}$ and conduct experiments for SONEX on Amazon dataset for GDRO task and ALEXR2 on Adult dataset for AUC maximization with ROC fairness constraints. The results are shown in table 3 which indicates the importance of setting $\beta\leq 1.0$ .
Results for varying $\theta$
We analyze the influence of different $\theta$ of ALEXR2 on model performance, testing $\theta\in\{\frac{1}{16},\frac{1}{8},\frac{1}{4},\frac{1}{2},1\}$ and conduct experiments for ALEXR2 on Adult dataset. As shown in Table 4, the final AUC of ALEXR2 remains relatively stable across this range, suggesting that it is not highly sensitive to $\theta$.
Benefit of Adam-type update
Finally we compare Adam-type update with momentum-type update. We conduct experiments for SONEX on Amazon dataset and ALEXR2 on Adult dataset and the results are summarized in table 5. They clearly demonstrate the advantage of using Adam-type updates, which consistently yield superior performance.
Table 3: Final loss and AUC of SONEX and ALEXR2 on Amazon dataset and Adult dataset respectively, with varying $\beta$
SONEX w/ varying $\beta$ 0.1 0.3 0.5 1.0
Final loss 0.5657 0.5563 0.5768 0.578
ALEXR2 w/ varying $\beta$ 0.1 0.3 0.5 1.0
Final AUC 0.8975 0.8976 0.8973 0.8969
Table 4: Final AUC of ALEXR2 with varying $\theta$ on Adult dataset
varying $\theta$ 0.1 0.3 0.5 1.0
Final AUC 0.8975 0.8976 0.8973 0.8969
Table 5: Comparison between Adam-type update and Momentum-type update for SONEX and ALEXR2 on Amazon dataset and Adult dataset, respectively
Algorithm Adam-type Momentum-type
SONEX(Final Loss) 0.5657 0.9666
ALEXR2(Final AUC) 0.8975 0.861
F.5 Verification of Assumption 4.3
@@ F.4 Ablation Study
F.6 Other Details of Experiments